What is the formula for electric potential at a point p on z axis from a system of 3 charges?

Grade:12

1 Answers

Ashish Kumar
7 years ago
let the 3 charges = q1,q2,q3. 
Let the distance of each charge from point = r1,r2,r3
Potential at p due to q1 = kq1/r1
Potential at p due to q2 = kq2/r2
Potential at p due to q3 = kq3/r3
Now potential is a scalar quantity so add all these with their signs normally.
Do approve the answer if u found it useful!